# Bootstrapping Object-level Planning with Large Language Models

### v1

@@ IV-A Object-level to Task-level Planning
The aim of task-level planning is to find a robot-executable task plan $\mathcal{P}_{\mu}$ solving a task $\mathcal{T}$. In more detail, a task plan is comprised of a sequence of smaller plan segments for each functional unit, i.e., $\mathcal{P}_{\mu}=\{\tilde{\mathcal{P}}_{\mu_{1}},\tilde{\mathcal{P}}_{\mu_{2}},...,\tilde{\mathcal{P}}_{\mu_{n}}\}$, where $\tilde{\mathcal{P}}_{\mu_{i}}=\{a_{\mu_{1}},a_{\mu_{2}},...,a_{\mu_{m}}\}$ denotes a plan segment achieving the subgoals described by a functional unit $\mathcal{FU}_{i}$ and $a_{\mu_{i}}$ refers to the $i$-th step corresponding to a parameterized skill in $\mathcal{A}$.
@@
PDDL solvers require two components: 1) a domain definition, and 2) a problem definition [10]. A domain definition provides details on what actions can be taken by a robot or agent as well as possible object types, while a problem definition captures the initial state of the robot and its environment ($s_{\mu}$) as well as the goal state ($g_{\mu}$) as logical predicates. In this work, we assume that we have a predefined domain definition with planning operators corresponding to a repertoire of parameterized, robot-executable skills $\mathcal{A}$. However, for problem definition generation, both $s_{\mu}$ and $g_{\mu}$ are adapted from a functional unit $\mathcal{FU}\in\mathcal{G}_{\mathcal{T}}$: predicates are constructed based on object-state pairs in $\mathcal{FU}$. In other words, transforming a FOON into PDDL requires the mapping of object properties defined by each object $o$ to predicates (where $o\in\mathcal{O}_{in}\cup\mathcal{O}_{out}$).
@@
Object-centered Predicates: We use object-centered predicates [24, 16], which describe constraints for collision-free action and motion. They are written as (⟨rel⟩ ⟨?obj_1⟩ ⟨?obj_2⟩), where ⟨rel⟩ refers to a geometric relation using the spatial adpositions in, on, or under, while ⟨?obj_1⟩ and ⟨?obj_2⟩ refer to objects described by a given predicate; These relations are described from the reference frame of each object, which permits propagating motion constraints at task planning for the generation of feasible plans [25]. For example, the predicate (on block1 block2) denotes that block2 is lying on top part of block1. We also use a virtual object air to describe free space in or on objects, which is important for collision-free picking (e.g., (on block1 air)—nothing is on a block, making it free for grasping).
Grounding: Each subgoal in an OLP (i.e., functional unit in FOON) must be grounded to the robot’s environment for effective task-level planning. For starters, object-level aliases must be linked to object references at the task level. This can be likened to how we as humans use recipes: a recipe refers to ingredients with words, but we must resolve their references to object instances around us when completing a recipe. This work assumes that there exists an exact mapping of objects described in an OLP to those existing in the environment, and we prompt the LLM to map each alias to an instance. For example, if we have two red blocks as objects in FOON, an LLM will map them to object instances red_block_1 and red_block_2 (Figure 4). Once these are resolved, we can then map object-state pairs described in FOON to task-level predicates: we use object poses (both position and orientation) and bounding boxes to derive object-centered predicates for each object $o$ in $\mathcal{G}_{\mathcal{T}}$ using the mechanism from prior work [25].
@@
Figure 4: Example of task-level grounding for an object-level plan (Figure 3). This state description is compatible with PDDL definitions (like Figure 5).
@@ IV-B Task-level to Motion-level Planning
With each plan segment $\tilde{\mathcal{P}_{\mu}}\in\mathcal{P}_{\mu}$, a robot can then execute a sequence of actions that will resolve object-level subgoals. We use motion-level planning to find collision-free robot movements that will resolve the effects of a robot’s skills. This work considers picking and placing actions (Figure 5). For the pick action (Figure 5a), we generate a trajectory that moves the robot’s end-effector from its initial position to a target object, while the place action (Figure 5b) moves the robot’s end-effector grasping an object from its initial pose to a position above a target surface or object. Initial and final poses of the hand for these actions can be directly obtained from object-centered hand-object relations encoded in the precondition and effects of their corresponding planning operators using rotation and translation geometric transformations [25].
@@
  ; collision-free constraints
  (in hand air)
  (on ?obj air)
  ; object is on surface
@@
  ; hand contains object
  (in hand ?obj)
  (not (in hand air))
  ; object is grasped
@@
  ; nothing is on surface
  (on ?surface air)
@@
  ; hand contains object
  (in hand ?obj)
  (on ?obj hand)
  ; collision-free constraints
@@
  (under ?obj air))
@@
  ; hand no longer contains object
  (in hand air)
  (not (in hand ?obj))
  ; nothing is on object
  (on ?obj air)
  (not (on ?obj hand))
  ; object is on surface
@@
Task Setting Planning Approach % Plan Complete % Success Avg. Plan Time (s) Avg. Tokens Avg. Plan Length
Tower OLP $\mathbf{86.00\%}$ $88.37\%$ $\mathbf{0.0043\pm 0.0021}$ $2406.38\pm 335.0091$ $10.2791\pm 4.5687$
LLM-Planner $44.00\%$ $77.27\%$ $12.0486\pm 6.3784$ $\mathbf{744.94\pm 120.9533}$ $10.2791\pm 4.5687$
LLM+P $18.00\%$ $\mathbf{94\%}$ $0.0346\pm 0.0312$ $1656.42\pm 170.2912$ $\mathbf{8.5556\pm 3.9291}$
DELTA $\mathbf{86.00\%}$ $69.77\%$ $0.0067\pm 0.01206$ $4871.88438.2610$ $9.0233\pm 4.7883$
Spelling OLP $\mathbf{80.00\%}$ $\mathbf{77.50\%}$ $0.02715\pm 0.0828$ $2588.66\pm 379.1376$ $\mathbf{8.45\pm 3.4932}$
LLM-Planner $22.00\%$ $44.44\%$ $7.5734\pm 2.4650$ $\mathbf{754.06\pm 102.7516}$ $9.7778\pm 3.3529$
@@
DELTA $78.00\%$ $64.10\%$ $\mathbf{0.0075\pm 0.0061}$ $4836.72\pm 475.0059$ $10.5641\pm 5.5998$
Organize OLP $\mathbf{81.43\%}$ $94.74\%$ $\mathbf{0.0080\pm 0.0053}$ $3051.5\pm 499.4818$ $15.3684\pm 7.1630$
LLM-Planner $35.71\%$ $64.00\%$ $24.1510\pm 15.9711$ $\mathbf{885.0571\pm 126.3086}$ $\mathbf{8.40\pm 2.2361}$
LLM+P $37.14\%$ $\mathbf{100.00\%}$ $0.0538\pm 0.1139$ $1891.3286\pm 212.4083$ $11.3077\pm 4.1547$
DELTA $67.14\%$ $80.85\%$ $0.0139\pm 0.0312$ $5329.9571\pm 470.9973$ $13.8298\pm 6.4042$
@@ V Evaluation
We evaluate the flexibility of our hierarchical planning approach, which exploits a LLM for extracting OLPs, which are transformed into task- and motion-level problems. Particularly, we highlight that a LLM cannot reliably produce PDDL definitions, and are unable to reliably task plan; however, we can prompt a LLM for object-level details useful to construct PDDL subgoals similar to prior work [16]. We compare our approach (OLP in Table I) with alternative methods on simulated experiments across several tasks.
@@ V-A Experimental Setup
We perform experiments in a simulated table-top environment in CoppeliaSim [26] with a Franka Emika Panda robot affixed to a table upon which blocks are randomly initialized. Given a task in natural language, the robot must perform a series of pick and place actions that will fulfill the task. We assume that the state of the environment is fully observable, where object poses and bounding boxes are known for perception. This information is used in motion-level planning to derive collision-free trajectories (if one exists for a given task plan action). In addition to Chat-GPT [^1] We tested gpt-4, gpt-4o, and chatgpt-4o-latest, but found chatgpt-4o-latest to produce better plans. See more about these model variants here: https://platform.openai.com/docs/models. [21] as our LLM of choice, we use Fast Downward [27], an off-the-shelf PDDL solver, for task-level planning in our method as well as baselines (discussed in Section V-B). When planning with Fast Downward, we use the A* algorithm with the landmark cut (LMCUT) heuristic for plan optimality. For collision-free motion-level planning, we use the RRT-Connect [28] algorithm as provided by OMPL [29] (included with CoppeliaSim).
Figure 6: Example of the initial and final states for the organizing table task. The tower building and spelling tasks involve stacking similar lettered blocks.
Task Settings: We design scenarios in which the robot has to complete several tasks for three tasks of increasing difficulty: 1) tower building, 2) spelling, and 3) organizing a table (Figure 6). The tower building task involves the robot assembling a tower of blocks of a given height $n$, where $3\leq n\leq 7$, with $n+1$ blocks provided on the table. The spelling task also involves robot constructing a tower of blocks of some height $n$, but with the added constraint that they correctly spell a given word of length $n$. This requires correct placement of lettered blocks, thus heavily depending on the LLM’s ability to generate the correct sequence of pick and place actions. Finally, the organizing involves a robot making piles of matching blocks: here, we initialize a scene of $3$ block types each with $n$ block instances (where $2\leq m\leq 4$). This can be seen as a mix of the two prior tasks, where alike but varying numbers of blocks must be placed into piles.
Metrics: We report the following metrics: 1) plan completion, which measures the percentage of all plans that were executed from start to finish; 2) success, which measures the percentage of all completely executed plans that were successful (i.e., if the final state matches the task prompt); 3) average plan computation time (in seconds); 4) average number of tokens for LLM prompting; and 5) average plan length across all successful executions.
@@ V-B Baseline Methods
We compare our OLP-based method to several baseline methods, for which we provide details below. These baseline methods also use Chat-GPT to either directly output a task plan or PDDL definitions, following the tracks of LLM-based planning work previously introduced.
@@ V-B1 LLM-Planner
This baseline centers the LLM as a planner, serving as a proxy for such methods [1, 2]. We directly provide a textual description of the state of the robot’s environment (denoted by $\tilde{s}$) and the robot’s executable skills ($\mathcal{A}$) as input and retrieve a task plan $\mathcal{P}_{\mu}$ as output. We then parse each action to identify target objects and surfaces needed for pick and place actions while performing the necessary motion-level planning to successfully resolve each action. This baseline approach evaluates the LLM’s ability to reason about the robot’s embodiment and produce a correct task plan.
@@ V-B2 LLM+P
LLM+P [6] uses a LLM to generate a PDDL problem definition given a text description of a task planning domain and initial state. As input to the LLM, we provide a description of the robot’s environment ($\tilde{s}$) and an example of a problem definition task, and we obtain a problem definition as output. We then use this output with our domain definition of predefined skills to acquire a task plan using Fast Downward [27], and this task plan is executed and resolved with motion-level planning. This baseline approach evaluates the LLM’s ability to accurately generate a PDDL problem file, compatible with a predefined set of skills, without explicitly performing object-level planning and reasoning.
@@ V-B3 DELTA
DELTA [8] is a task planning method which auto-regressively prompts a LLM to derive PDDL domain and problem definitions. Similar to our approach, a task $\mathcal{T}$ is broken down into sub-goals, each of which are formulated as their own sub-goal PDDL problem file. We prompt the LLM with details about robot actions ($\mathcal{A}$) as well as the objects available to the robot, after which a domain file is generated. The LLM is then provided with state description $\tilde{s}$ and the task prompt $\mathcal{T}$ to generate a problem file containing all goals (similar to the output of LLM+P [6]). This problem file is then broken down into sub-goal problem files based on PDDL sub-goals auto-regressively suggested by the LLM; this scopes the problem into sub-goal actions that are akin to functional units. We hypothesize that although this method will create simpler and smaller problem definitions, this heavily relies on the LLM’s ability to generate syntactically and semantically correct definitions, which may not be as reliable as our method.
@@ V-C Results and Discussion
Our experimental results show that our OLP-based method performs better than baselines that either generate a task plan or PDDL files (Table I). Across all tasks and approaches, plans were not completely executable due to motion-level planning failures, where plans were not found in a reasonable amount of time. Despite this phenomenon, our OLP approach produces the most plan completions across all task settings on average (Figure 7). While not always successful in execution, our OLP approach, generates plans that exhibit high success rate, matching the intention of the given instruction. Interestingly, the spelling task showed the lowest success rate across all approaches: we attribute this to incorrect reasoning done by the LLM at both object and task levels.
@@
Although LLM-Planner generates task plans without a solver, it does not complete a majority of tasks because the LLM has poor understanding of the configuration of the robot’s environment for collision-free motion. Failure cases involve attempts to pick up an object blocked by another object or place an object on an occupied spot. LLM+P also exhibits poor performance: although the LLM is capable of directly outputting PDDL, failures were mainly attributed to inaccurate problem definitions. This may be attributed to the fact that LLM+P uses fewer prompts than OLP and DELTA; also, unlike DELTA, LLM+P does not give definitions of PDDL planning operators, thus providing less context to the LLM. We also observed that PDDL problems generated by LLM+P and DELTA were susceptible to incorrect syntax, which is a drawback of LLM-based PDDL generation. DELTA, whose approach closely resembles our method, performs better than LLM+P and LLM-Planner baselines, but it does not perform as well as our method while also generally requiring more tokens on average to generate planning definitions. Similar to DELTA, OLP also demonstrates the advantage of bootstrapping task-level planning with PDDL subgoal definitions (reflected by low average planning times) but without relying upon the LLM to correctly generate PDDL definitions. Our approach also requires less interaction with the LLM than DELTA as reflected by the average number of tokens.
@@
Figure 7: Graph showing percentage of plans completely executed across all tasks and approaches for different number of blocks in the scene.
@@
Limitations: Much like how we humans plan, object-level planning serves as a critical interface between language and TAMP. However, we acknowledge several shortcomings of our work: 1) our approach requires explicit robot skill definitions in PDDL, which may not always transfer across robot systems; 2) we assume a set of FOON samples for few-shot learning; 3) we did not consider plan recovery if objects were knocked down during execution (thus potentially dropping the success rate of completely executed plans); and, most importantly, 4) as with baselines, this approach depends on a correctly generated object-level plan compatible with task-level planning for subgoal definitions. Our evaluations are solely done for pick-and-place tasks, which do not highlight the benefits of the semantic richness of object-level plans. As future work, we will explore broader task diversity and adapting plans to novel settings like prior work [30, 31]. Like recent work [32], we can also integrate human feedback to correct LLM-generated errors at the object level. Further, we will explore learning from demonstration to acquire task-level domain definitions to address this work’s assumption of predefined skills.
VI Related Work
Language Models for Planning: Many researchers have explored the use of language models for robotics applications, having been inspired by their remarkable performance in language-related tasks. Prior work have investigated the planning capabilities of LLMs [33, 12]. Other works supplement task planning with language models [6, 34, 5, 35, 8, 32]. LLM+P [6] generates PDDL problem file via LLM prompting. Much like our work, recent work use LLMs as an informer of subgoals for classical planning [35, 8, 32] In particular, DELTA [8] resembles our method in that it decomposes a task into a series of PDDL subgoal definitions directly outputted by a LLM. Our approach uses a LLM at the object level and not task level (i.e., PDDL). Recent work also iteratively prompts a LLM for generating FOONs [36]. Similar to DELTA, they do not focus on generating nor executing physically valid plans.
Language Models as Planners: Several works treat language models as robotic task planners. SayCan [1] combines a language model and affordance detectors for driving robotic execution given a task prompt. PaLM-E [2] is an embodied language model that directly incorporates continuous observations (like images, state estimates, or other sensor modalities) into the language embedding space. These works have shown that LLMs are capable of performing some degree of embodied reasoning. However, one major drawback to these works is that they require a large amount of engineering effort, particularly for them to operate in novel environments or to perform long-horizon tasks. Prior work also exploit the reasoning capabilities of a LLM for solving a wide range of tasks both in simulation [3, 37, 4] and with a real robot [4].
VII Conclusion

### v3

@@ IV-A Object-Level to Task-Level Planning
The aim of task-level planning is to find a robot-executable task plan $\mathcal{P}_{\mu}$ that solves task $\mathcal{T}$. A task plan is composed of a sequence of smaller plan segments for each functional unit, i.e., $\mathcal{P}_{\mu}=\{\tilde{\mathcal{P}}_{\mu_{1}},\tilde{\mathcal{P}}_{\mu_{2}},...,\tilde{\mathcal{P}}_{\mu_{n}}\}$, where $\tilde{\mathcal{P}}_{\mu_{i}}=\{a_{\mu_{1}},a_{\mu_{2}},...,a_{\mu_{m}}\}$ denotes a plan segment achieving the subgoals described by a functional unit $\mathcal{FU}_{i}$ and $a_{\mu_{i}}$ refers to the $i$-th step corresponding to a parameterized skill in $\mathcal{A}$.
@@
PDDL solvers require two components: a domain definition and a problem definition [11]. A domain definition provides details on what actions can be taken by a robot as well as possible object types, while a problem definition captures the initial state of the robot and its environment ($s_{\mu}$) as well as the target goal state ($g_{\mu}$) as logical predicates. We assume a predefined domain definition with planning operators corresponding to a repertoire of parameterized, robot-executable skills $\mathcal{A}$. For the generation of problem definitions, both $s_{\mu}$ and $g_{\mu}$ are adapted from a functional unit $\mathcal{FU}\in\mathcal{G}_{\mathcal{T}}$: predicates are constructed based on the object-state pairs in $\mathcal{FU}$ [17]. In other words, transforming a FOON into PDDL requires mapping attributes of each object $o$ to predicates (where $o\in\mathcal{O}_{in}\cup\mathcal{O}_{out}$).
@@
Object-centered Predicates: We use object-centered predicates [33, 17] that describe constraints for collision-free motion. They are written as (⟨rel⟩ ⟨?obj_1⟩ ⟨?obj_2⟩), where ⟨rel⟩ refers to a geometric relation using the spatial adpositions in, on, or under, while ⟨?obj_1⟩ and ⟨?obj_2⟩ refer to objects described by a given predicate. These relations are described from the reference frame of each object, which permits propagating motion constraints at task planning for the generation of feasible plans [34]. For example, the predicate (on block_1 block_2) denotes that block_2 lies on top of block_1. We also use a virtual object air to describe free space in or on objects, which is important for collision-free picking, i.e., (on block_1 air)—nothing is on a block, which makes it free for grasping).
@@
Figure 4: Example of task-level grounding for an object-level plan (Figure 3), which is compatible by design with the planning operators in Figure 5.
Grounding: Each subgoal in an OLP (i.e., functional unit in FOON) must be grounded to the robot’s environment for effective task-level planning. For starters, object-level aliases must be linked to object references at the task level. This can be likened to how we as humans use recipes: a recipe refers to ingredients with words, but we must resolve their references to object instances around us when completing a recipe. This work assumes that there exists an exact mapping of objects described in an OLP to those existing in the environment, and we prompt the LLM to map each alias to an instance. For example, if we have two red blocks as objects in an OLP (Figure 3), an LLM will map them to object instances red_block_1 and red_block_2 (Figure 4). Once completed, we obtain a mapping of object-state pairs to task-level predicates: we use object poses (both position and orientation) and bounding boxes to derive object-centered predicates for each object $o$ in $\mathcal{G}_{\mathcal{T}}$ using the mechanism from previous work [34].
@@ IV-B Task-Level to Motion-Level Planning
With each plan segment $\tilde{\mathcal{P}_{\mu}}\in\mathcal{P}_{\mu}$, a robot can then execute a sequence of actions that resolve object-level subgoals. We use motion-level planning to find collision-free robot movements that will achieve the effects of a robot’s skills. This work considers picking and placing actions (Figure 5). For the pick action (Figure 5a), we generate a trajectory that moves the robot’s end-effector from its initial position to a target object, while the place action (Figure 5b) moves the robot’s end-effector grasping an object from its initial pose to a position above a target surface or object. The initial and final poses of the hand for these actions can be obtained directly from object-centered hand-object relations encoded in the preconditions and effects of their corresponding planning operators using geometric rotation and translation transformations [34].
@@
  ; collision-free constraints:
  (in hand air) (on ?obj air)
  ; object is on a surface:
@@
  ; hand contains target object:
  (in hand ?obj) (not (in hand air))
  ; object has been grasped:
@@
  ; nothing is on surface:
   (not (on ?surface ?obj))
@@
  ; collision-free constraints:
@@
  (under ?obj air)
  ; hand contains object:
  (in hand ?obj) (on ?obj hand) )
@@
  ; hand no longer contains object:
  (in hand air) (not (in hand ?obj))
  ; object is on surface:
@@
Task Setting Planning Approach % Plan Complete $\uparrow$ % Success $\uparrow$ Avg. Plan Time (s) $\downarrow$ Avg. Tokens $\downarrow$ Avg. Plan Length $\downarrow$
Tower OLP $\mathbf{86.00\%}$ $\mathbf{76.00\%}$ $\mathbf{0.0043\pm 0.0021}$ $2406.38\pm 335.0091$ $10.2791\pm 4.5687$
LLM-Planner $44.00\%$ $34.00\%$ $12.0486\pm 6.3784$ $\mathbf{744.94\pm 120.9533}$ $10.2791\pm 4.5687$
LLM+P $18.00\%$ $34.00\%$ $0.0346\pm 0.0312$ $1656.42\pm 170.2912$ $\mathbf{8.5556\pm 3.9291}$
DELTA $\mathbf{86.00\%}$ $60.00\%$ $0.0067\pm 0.01206$ $4871.88438.2610$ $9.0233\pm 4.7883$
Spelling OLP $\mathbf{80.00\%}$ $\mathbf{62.00\%}$ $0.02715\pm 0.0828$ $2588.66\pm 379.1376$ $\mathbf{8.45\pm 3.4932}$
LLM-Planner $22.00\%$ $16.00\%$ $7.5734\pm 2.4650$ $\mathbf{754.06\pm 102.7516}$ $9.7778\pm 3.3529$
@@
DELTA $78.00\%$ $50.00\%$ $\mathbf{0.0075\pm 0.0061}$ $4836.72\pm 475.0059$ $10.5641\pm 5.5998$
Organize OLP $\mathbf{81.43\%}$ $\mathbf{77.14\%}$ $\mathbf{0.0080\pm 0.0053}$ $3051.5\pm 499.4818$ $15.3684\pm 7.1630$
LLM-Planner $35.71\%$ $22.86\%$ $24.1510\pm 15.9711$ $\mathbf{885.0571\pm 126.3086}$ $\mathbf{8.40\pm 2.2361}$
LLM+P $37.14\%$ $37.14\%$ $0.0538\pm 0.1139$ $1891.3286\pm 212.4083$ $11.3077\pm 4.1547$
DELTA $67.14\%$ $54.29\%$ $0.0139\pm 0.0312$ $5329.9571\pm 470.9973$ $13.8298\pm 6.4042$
@@ V Evaluation
We evaluate the flexibility of our approach (denoted as OLP in Table I) with alternative methods on several tasks in simulated experiments. Our results show that an LLM cannot reliably produce PDDL definitions and is unable to reliably task plan due to its lack of spatial understanding; however, we can prompt an LLM for object-level subgoals compatible with our modular approach from previous work [17].
@@ V-A Experimental Setup
We perform experiments in a simulated table-top environment in CoppeliaSim [35] with a Franka Emika Panda robot affixed to a table upon which blocks are randomly initialized. Given a task specified in natural language, the robot must perform a sequence of pick and place actions (defined in Figure 5) fulfilling the task. We assume that the state of the environment is fully observable—object poses and bounding boxes are known via perception. This information is used in motion-level planning to generate collision-free trajectories. In addition to Chat-GPT [^3] We tested gpt-4, gpt-4o, and chatgpt-4o-latest, but found chatgpt-4o-latest to produce the best plans, adhering to instructions. [22] as our LLM of choice, we use Fast Downward [36], an off-the-shelf PDDL solver, for task-level planning in our method and baselines (discussed in Section V-B). When planning with Fast Downward, we use the A* algorithm with the landmark cut (LMCUT) heuristic for plan optimality. For motion-level planning, we use RRT-Connect [37] as provided by OMPL [23].
Task Settings: We design scenarios in which the robot has to complete several tasks for three tasks of increasing difficulty: 1) tower building, 2) spelling, and 3) organizing a table (Figure 6). The tower building task involves the robot assembling a tower of blocks of a given height $n$, where $3\leq n\leq 7$, with $n+1$ blocks provided on the table. The spelling task also involves robot constructing a tower of blocks of some height $n$, but with the added constraint that the blocks correctly spell a given word of length $n$. This requires correct placement of lettered blocks, thus heavily depending on the LLM’s ability to generate the correct sequence of pick and place actions. Finally, the organizing task involves a robot making piles of matching blocks: here, we initialize a scene of $3$ block types, each with $n$ block instances (where $2\leq m\leq 4$). This can be seen as a mix of the two prior tasks, where alike but varying numbers of blocks must be placed into piles.
Metrics: We report the following metrics: 1) plan completion, which measures the percentage of all plans that were executed from start to finish regardless of whether the task objective was achieved; 2) success, which measures the percentage of successfully executed plans that achieve the task objective; 3) average plan computation time (in seconds); 4) average number of tokens for LLM prompting; and 5) average plan length across all successful executions.
Figure 6: Example of initial and final states for the spelling and organizing table tasks. The tower building task is akin to spelling without ordering constraints.
@@ V-B Baseline Methods
We compare our OLP-based method to several baseline methods, for which we provide details below. These baseline methods also rely upon Chat-GPT to either directly output a task plan or PDDL definitions, following the tracks of LLM-based planning work previously introduced.
@@ V-B1 LLM-Planner
This baseline serves as a proxy for methods that directly plan with LLMs [1, 2]. We directly provide a textual description of the state of the robot’s environment (denoted by $\tilde{s}$) and the robot’s executable skills ($\mathcal{A}$) as input and retrieve a task plan $\mathcal{P}_{\mu}$ as output. We then parse each action to identify target objects and surfaces needed for pick and place actions while performing the necessary motion-level planning to successfully resolve each action. This baseline approach evaluates the LLM’s ability to reason about the robot’s embodiment and produce a correct task plan.
@@ V-B2 LLM+P
LLM+P [6] uses an LLM to generate a PDDL problem definition given a text description of a task planning domain and the initial state of the scene. As input to the LLM, we provide a description of the robot’s environment ($\tilde{s}$) and an example of a problem definition task, and we obtain a problem definition for task $\mathcal{T}$ as output. We then use this output with our domain definition of predefined skills to acquire a task plan using Fast Downward [36], and this task plan is executed and resolved with motion-level planning. This baseline approach evaluates the LLM’s ability to accurately generate a PDDL problem file, compatible with a predefined set of skills, without explicitly performing object-level planning and reasoning.
@@ V-B3 DELTA
DELTA [8] is a task planning method that auto-regressively prompts an LLM to derive PDDL domain and problem definitions. Similar to our approach, a task $\mathcal{T}$ is broken down into subgoals, each of which is formulated as their own subgoal PDDL problem file. We prompt the LLM with details about robot actions ($\mathcal{A}$) as well as the objects available to the robot, after which a domain file is generated. The LLM is then provided with a state description $\tilde{s}$ and a task prompt $\mathcal{T}$ to generate a problem file that contains all goals (similar to the output of LLM+P [6]). This problem file is then broken down into subgoal problem files based on PDDL subgoals auto-regressively suggested by the LLM; this scopes the problem into subgoal actions that are akin to functional units. We hypothesize that although this method will create simpler and smaller problem definitions, it heavily relies on the LLM’s ability to generate syntactically and semantically correct definitions, which may not be as reliable as our method.
@@ V-C Results and Discussion
Our experimental results show that our OLP-based method performs better than baselines that either directly generate a task plan or PDDL files (Table I). Across all tasks and evaluated approaches, we found that some plans were not fully executable due to motion-level planning failures, where the plans were not found in reasonable time. Despite this phenomenon, our approach produces the most plan completions in all task settings on average (Figure 7). Although OLP was not always successful in execution, our approach generates plans that exhibit the highest success rates, matching the intention of the given instruction. Interestingly, the spelling task showed the lowest success rate in all approaches. We attribute this to incorrect reasoning performed by the LLM at both the object and task levels, where the LLM may generate a plan sketch to stack the blocks in an incorrect or reversed order.
@@
Although LLM-Planner generates plans without a solver, it does not complete a majority of tasks because the LLM poorly understands the configuration of the robot’s environment for collision-free motion. As a result, it incorrectly proposes actions that attempt to pick up an object blocked by another object or place an object in an occupied spot. LLM+P also exhibits poor performance: although the LLM is capable of directly outputting PDDL, failures were mainly attributed to inaccurate problem definitions. This may be due to the fact that LLM+P uses fewer prompts than OLP and DELTA; also, unlike DELTA, LLM+P does not provide definitions of PDDL planning operators, thus providing less context to the LLM. We also observed that the PDDL problems generated by LLM+P and DELTA were susceptible to incorrect syntax, which is a drawback of LLM-based PDDL generation. DELTA, whose approach closely resembles our method, performs better than LLM+P and LLM-Planner baselines, but it does not perform as well as our method while also generally requiring more tokens on average to generate planning definitions. Similar to DELTA, OLP also demonstrates the advantage of bootstrapping task-level planning with PDDL subgoal definitions (reflected by low average planning times) but without relying upon the LLM to correctly generate PDDL definitions. Our approach also requires less interaction with the LLM than DELTA as reflected by the average number of tokens.
@@
Figure 7: Graph showing percentage of plans completely executed using all approaches for different number of blocks across tasks (best viewed in colour).
@@
Limitations: Much like how we humans plan, object-level planning serves as a critical interface between language and TAMP. Our approach requires robot skill definitions specified as PDDL, which may not always transfer across robot systems. However, we assume a set of FOON samples for few-shot learning. Further, we did not consider plan recovery if objects were knocked down during execution, thus lowering the success rate of completely executed plans. Most importantly, as with baselines, this approach depends on a correctly generated object-level plan compatible with task-level planning for subgoal definitions. In addition, our evaluations are performed solely on pick-and-place tasks, which do not highlight the benefits of the semantic richness of object-level plans. In future work, we will explore broader task diversity and examine how we can use an LLM to adapt existing object-level plans to novel scenarios similar to prior work [38, 39]. Like recent work [27], we can also integrate human feedback to correct LLM-generated errors at the object level. We will also explore learning from demonstration to acquire task-level domain definitions to address our assumption of predefined skills.
VI Conclusion